Semans is a manufacturer that produces bracket assemblies. Demand for bracket assemblies (X) is 128 units. The following is the BOM in indented form:

ITEM DESCRIPTION USAGE
X Bracket assembly 1
A Wall board 3
B Hanger subassembly 2
D Hanger casting 3
E Ceramic knob 2
C Rivet head screw 4
F Metal tong 4
G Plastic cap 2

Below is a table indicating current inventory levels:


Item X A B C D E F G
Inventory 30 21 62 24 189 270 1,200 105


b. What are the net requirements for each item? (Leave no cells blank - be certain to enter "0" wherever required.)


Item Net Requirements
X
A
B
C
D
E
F
G
Answer #1

Based on the given data, we prepare a product structure tree as shown below:

Bracket assembly (X) Wall Board (A) Hanger Subassembly (B) Rivet head screw Hanger casting (D) Ceramic Knob (E) Metal Tong (F

Based on the usage given and current inventory levels, we find the requirement of each part as shown below:

128 - 30 = 98 Bracket assembly (X) (1) 98*3-21 = 273 Wall Board (A) (3) 98*2- 62 = 134 Hanger Subassembly (B) (2) 98*4 - 24 =

The on-hand inventory is subtracted as shown in red. Hence,

Item Net Requirements
X 98
A 273
B 134
C 368
D 213
E 0
F 272
G - 631

-----------------------------------------------------------------------------------------------------------------------

In case of any doubt, please ask through the comment section before Upvote/downvote.

If you liked the answer, please give an upvote. This will be quite encouraging for me, thank-you!!!